m К . Х3
druns armas on de System co hili la X, and to x = 183 1x d X and X2

make the transformation of the system to the original coordinate
Answer #1

3) mK ht I mm * - on inson som monit egn of motion of the first mass: = - K + KLX-4) - K KK - - (2-4) = -2KG + KA2 + 2K - KR2o wint 2K -mwnt2K -K K -mw. M2+2K ng then if we want to find normal which are non-trivial (ie 4 + 8 -Wm+26 - 0 - mu+2K k = 0so, eqh reducep to. Mlay = -K 127 = -127 where 127= [ ] → eigen barip on normel modes. 3 mox -- ** ; maintenan; max-ax The niV12 12 V / - V ve p S o - 13/2 12 12/2 1/2) vuz in diagonal basis, 12=2K) = 19 = (2+J£)K), - 17 = (2-12) k) d - - in original